Integrate each function without using the table. a) ſcosx In(sinx) dx b) si V x² -16 •doc +2
Answer #1

s log (sinx) (0 sac, doo ret t=Sina → dt = Os ac doc s log (tidt Slog (sinx) (050dx = using integrations by part & udv. uv-su5 (6) Jac²_16 dac 2 4 26 solve √2²-16 dae 202 using Integration by parte udv uv-sudu u=52²-16 , v= 1 cz w I apo v=-1 5x²-16 dAnd 1522-16 dx = 26 + log etn²- then +C 6 dt - 15 민 16 1 t 조 4 C 4 5 (log |a+Jn16lo 3 5 (룹 +) + iogg - log+ 읗t 10922- 1092 읗